OpenStudy (anonymous):

Factor completely (x+5)^2 - 2(x+5)(x-5) + (x-5)^2

OpenStudy (anonymous):

replace (x+5) with u, and replace (x-5) with v, and it might be easier to factor.

OpenStudy (anonymous):

replace (x+5) with u, and replace (x-5) with v, and it might be easier to factor.

OpenStudy (anonymous):

and then change them back to u = (x+5) and v = (x - 5)

OpenStudy (anonymous):